Hello I need help with number # 2. as it’s asking me to make predictions and fill in the first column of Part A of this experiment
d. 2.5 xIO 10.t 2. Make predictions and fill in the first column of Part A of this experiment 3. What is the app roximate expected pH ofa 0.030 M HNo, solution? hat is the final pH after 1 drop (0.05 mL) of 6 M HCl is added to 1.01 freshly prepared pure wat a significant pH change? 0.05ml (demı) ( er that was originally at a pH of 7.0. Is there ): .000 3 mol H
media%2Ff0e%2Ff0ef7976-a327-4dfa-8ee6-29
Please help as I don’t know what problem two means by predicting for the first column?
Answer #1

2).ans:- We know that, pH of a solution is the Negative logarithm of hydronium ion or hydrogen ion concentration.i.e.,

pH = -log [H3O+]or -log [H+]

0.1 M HCl = 10-1 M HCl ------> H+ + Cl-

here, [H+] = 10-1 M , Therefore, pH = -log [10-1] = - (-1) = 1

0.01 M HCl =  10-2 M HCl ------> H+ + Cl-

here ,[H+] = 10-2 M Therefore, pH = -log [10-2] = - (-2) = 2

0.001 M HCl = 10-3 M HCl ------> H+ + Cl-

here ,[H+] = 10-3 M Therefore, pH = -log [10-3] = - (-3) = 3

0.1 M HC2H3O2or  10--1 M CH3COOH(Acetic acid)   ------> CH3COO- + H+

Acetic acid is weak acid therefore , It does not dissociate completely in aqueous solution and thus, pH is not 1 as for 0.1 M HCl. In such cases for weak acids or bases Ostwald’s Dilution Law is used .i.e.,

pH = - log[H+] = -log Cα and where, α=√ (Ka / C)

Where , C= concentration, Ka = Dissociation Consant , α = Degree of dissociation

Ka for acetic Acid is given as = 1.8 x 10-5

Thus,   α =  √ (Ka / C) = √ 1.8 x 10-5 / 0.1 = 1.34 x 10-2

pH = -log ( 0.1 X 1.34 x 10-2 ) = 2.87

Thus, pH of 0.1 M of HC2H3O2 is 2.87.

0.1 M NH4Cl It is a salt made from NH3 and HCl. It is thus, a salt of weak base and strong acid. pH for such salt is given by the following equation,

pH = 7 - 1/2 pKb -1/2 log C

pKb for NH4Cl is = 4.74 (pKb = -log Kb) and C = 0.1 M

Putting the values in above equation and pH obtained for 0.1 M NH4Cl is = 5.13

pH of Deionised water = Neutral i.e., 7

pH of Tap water = it is generally less than the neutral i.e. less than the 7 . or we can say somewhere between 6 and 7.

pH of 0.1 M NaCl = It is a neutral salt as it is formed of srong acid HCl and strong base NaOH. THus pH of 0.1 M NaCl is 7.

pH of 0.1 M NaC2H3O2= It is a salt of weak acid CH3COOH and strong Base NaOH.pH for this salt is given as,

pH = 7 + 1/2 pKa + 1/2 log C

Thus, pH = 5.13

pH of 0.1 M NaHCO3 = 8.3

pH of 0.1 M Na2CO3 = It is an basic salt hence pH is 11.6

pH of 0.1 M NH3 = It is moderate Base = approximately 12

pH of 0.1 M NaOH = 13 as NaOH------> Na+ + OH- .it is a strong base and produce OH- ions in the solution .

By knowing the values of Ka and Kb we can easily predict the pH values.